# A novel hybrid time-varying graph neural network for traffic flow forecasting

### v1

@@ IV-B1 Graph learning
In order to capture the static spatial correlations between traffic nodes, we use two learnable embedding vectors to adaptively learn the static adjacency matrix at each time step. The static adjacency matrix of the time step is defined as follows:
where is a trainable node embedding, is a trainable bias of the node embedding . Herein, in order to ensure the diversity of at different time step, we set different bias for different time step during the training process, Relu() is used as the activation function.
In order to model the similarity of traffic flow between different traffic nodes, a traffic pattern graph is defined by calculating the correlations of historical traffic flows using the DTW algorithm [28]. In general, traffic pattern graph reflects a long-term traffic pattern. However, short-term traffic pattern may play an more important role in short-term traffic flow prediction. Therefore, in order to obtain more spatial topology-related information at different time step, we have proposed a static graph learning module to improve the learning ability of the static adjacency matrix. At each time step, we can construct a static graph with the static adjacency matrix defined in Eq.(9). Then, in order to more accurately model static spatial correlations between different traffic nodes, we can further calculate the coupling time-varing adjacency matrix by the following Eq.(10):
where are learnable parameter matrices.after completing the graph coupling process, we substitute the time-varying graph in at each corresponding time step with the coupled graph as the final .
@@
where is the value of the hidden embedding at time step . is the mapped traffic features including short-term dynamic spatial traffic patterns implicit in current traffic features . is the -th node in . represents a splicing operation. Indicates the dynamic traffic similarity between node and node in the road network.
Then, using Softmax() for normalization, the generation process of is as Eq. (13):
Then,we think dynamic changes of spatial correlations occuring in the current time should belong to the most active nodes. such as nodes which lied in adjency matrix and DTW matrix.thus,we combine the adjency matrix and DTW matrix as the mask matrix of dynamic spatial correlations.Finally the dynamic adjacency matrix can be define as Eq. (14):
IV-B2 Time-varying graph convolution gated recurrent unit
GRU is a variant of RNN which maintains the effectiveness of RNN in capturing temporal features while simplifying the parameters of the model, making it widely used in capturing temporal correlation. Draw inspiration from previous studies [29] [30], we modify the MLP layer in GRU with the proposed TVGCN model , and combine TVGCN and GRU to capture Spatio- temporal dependencies between different nodes. The specific form can be defined as Eq. (15):
where and are the input hidden emebedding and output hidden embedding at time step. represents the concatenation operation. and represent reset gate and update gate at time steps respectively. and represent the static graph operation and dynamic graph operation severally., and , and , are learnable parameters in TVGCRM
@@ IV-B2 Time-varying graph convolution gated recurrent unit
IV-C Position Encodings
Among them, represents the position and represents the feature dimension.
@@ V-A Datasets
We validate the predictive performance of HTVGNN by conducting experiments on four real-world traffic datasets, namely PeMSD3, PeMSD4, PeMSD7, and PeMSD8. These datasets were collected by the California Department of Motor Transportation Performance Measurement System (PeMS) [39]. Please refer to Table 1 for detailed information.
@@ V-B Evaluation Metrics
In this study, we employed three effectiveness measures, namely Mean Absolute Error (MAE), Mean Absolute Percentage Error (MAPE), and Root Mean Square Error (RMSE), to assess the accuracy of the HTVGNN model as described in Eq. 20 through 22.
@@ V-C Baselines
We compare the HTVGNN model against fourteen state-of-the-art baselines as outlined below:
@@
[33] ASTGCN introduces attention mechanisms into CNN and GCN to respectively model spatial and temporal dynamics of traf?c data.
[29] AGCRN utilizes a node adaptive parameter learning to enhance GCN and combines it with GRU to capture spatial and temporal features of traf?c data.
[34] use a spatial-temporal fusion module that a new graph structure constructs based on a data- driven.
[35] which designs an encode-decoder architecture to consistently model the spatial and temporal dependencies in the traffic time-series by GCN and transformer, respectively approach Dual Dynamic Spatial-Temporal Graph Convolution Network for Traffic Prediction.
@@
[37] The -Shape algorithm is used to classify regional nodes to model the delay between different nodes, thereby constructing a short-range graph mask matrix. In addition, DTW is used to construct another long-range graph mask matrix. On this basis, a split spatio-temporal multi-head attention mechanism is designed to model short-range spatial dependence, long-range spatial dependence and time dependence respectively.
[38] The encoder-decoder structure consists of a temporal Transformer and a spatial Transformer. The difference is that a recall module is added to the decoder to provide supplementary information for the prediction task. The designed DHM module is used to dynamically adjust the complexity of the model according to the complexity of the prediction task.
@@ V-D1 Dataset Preprocessing
Following recent methodologies [40], [41] and [42], we partition the training, validation, and test sets in a 6:2:2 ratio. To ensure more stable training, we employ z-score standardization. Additionally, we utilize historical traffic flow speed data from the past time slices (60 min) to predict future traffic flow speed in time slices (60 min). The input data is normalized using Z-Score as shown in Eq. (17).
@@ V-D2 Model Settings
The experiments were conducted on a machine equipped with dual NVIDIA RTX and 32GB of memory. HTVGNN was implemented using Ubuntu 18.04, PyTorch 2.0.0, and Python 3.9.7, with a hidden dimension E of 64, eight heads for the Transformer, and two layers for GRU. The optimal model was determined based on performance on the validation set, trained using Adam optimizer [43] with a learning rate of 0.001 over a batch size of 32 for training epoch of up to 300.
@@ V-E Performance Comparison
Tables I and II present the comparison results of 12 prediction models on four datasets. Overall, our HTVGNN achieved exceptional prediction accuracy across all 12 prediction strides. The deep learning-based approaches outperformed traditional statistical-based approaches (VAR), indicating their efficacy in modeling highly nonlinear traffic flows. As depicted in Tables I and II, these approaches exhibited suboptimal prediction performance. Moreover, methods that capture spatiotemporal dependencies generally surpassed those solely modeling temporal dependencies (LSTM), underscoring the significance of spatial dependencies. DCRNN and STGCN employ predefined graphs to model spatial dependencies in road networks; however, their performance heavily relies on the quality of these predefined graphs. AGCRN utilizes adaptive graphs to capture spatial dependencies, which typically enhances predictions compared to predefined graphs. Nevertheless, both predefined and adaptive graphs remain static during the prediction stage, disregarding dynamic changes in the road network. In contrast, models that dynamically model spatial dependencies (GMAN, PDFormer) significantly outperform other approaches based on static graph structures (DCRNN, STGCN, GWN). This can be attributed to their ability to learn dynamically changing graph structures and provide a broader representation space for modeling more complex spatiotemporal dependencies.Compared to other models, HTVGNN incorporates a time-varying mask matrix to enhance the multi-head attention mechanism for accurately capturing dynamic temporal correlations. Moreover, HTVGNN employs a dedicated coupled graph learning method tailored for multi-head attention, effectively improving its performance in long-term prediction tasks. Additionally, to better model dynamic spatial correlations in road networks, we define a mask based on topological and semantic matrices to optimize the dynamic time-varying graphs. Overall, our model outperforms existing state-of-the-art approaches on four datasets, achieving superior prediction accuracy.
Additionally, in order to demonstrate the overall superiority of our model, we provide visual representations for each prediction point of HTVGNN and 11 comparative models as depicted in Fig. 4. In terms of the PEMS03 and PEMS04 datasets, our model exhibits the smallest cumulative error as the time steps increase, highlighting a distinct advantage of HTVGNN during long-term prediction stages. Furthermore, for PEMS07 and PEMS08 datasets, our model consistently maintains a significant advantage across all time steps.
V-F Abliation sudy
@@
w/o ETPMSA: TVGCRM should be utilized exclusively, with the exclusion of the enhanced temporal perception multi-head self-attention mechanism. This variant primarily demonstrates the function of the enhanced temporal perception multi-head self-attention.
@@
w/o TV: Replace the time-varying graph in TVGCRM with a topological graph, emphasizing its pivotal role in capturing temporal dynamic correlations.
@@
w/o TR: Replace TVGCRM with a topological graph-based graph convolution, which primarily emphasizes the role of the RNN module in HTVGNN. At this juncture, the number of layers remains unchanged.
@@
The experiments were conducted multiple times, and the optimal MAE, RMSE, and MAPE values for each variant on PEMSD4 and PEMSD8 are presented in Table 4.
@@
The results presented in Table 4 demonstrate the superior performance of HTVGNN without ETPMSA, highlighting the crucial role played by ETPMSA in HTVGNN. Furthermore, the inferior performance of w/o TV compared to HTVGNN indicates that capturing dynamic correlations through time-varying graphs is essential. Additionally, w/o TR being inferior to HTVGNN reveals the positive contribution of RNN in capturing temporal correlations.
@@
Given that our model is a hybrid time-varying graph convolutional neural network, this tutorial aims to dissect the impact of time-varying embedding and time-varying mask embedding by dividing the model into modules, ultimately analyzing their influence on the model’s performance.
@@
V-F1 Effect analysis of time-varying embedding in RTP multi-head attention
We compare transformer modules based on enhanced temporal perception multi-head self-attention across different layers to analyze the performance improvement of the enhanced temporal perception multi-head self-attention Transformer, which utilizes an time-varing mask matrix, relative to the Transformer with temporal multi-head self-attention. Internally, our Transformer employs dynamic graph convolution based on topological graphs to capture dynamic spatial correlations. Prior to that, we investigate the impact of varying sizes of mask embedding and dropout on enhanced temporal perception multi-head self-attention. For this analysis, we set the number of layers to 1 and select optimal spatial embedding and dropout ratios. Fig. 6 illustrates how different time-varing mask embeddings and dropouts affect PEMS04 and PEMS08 datasets in terms of enhanced temporal perception multi-head self-attention.
@@ V-F1 Effect analysis of time-varying embedding in RTP multi-head attention
To highlight the advantages of enhanced temporal perception in multi-head self-attention, we set the number of layers to 4 based on optimal mask embedding and dropout. We designed two variants as described below:
@@
| PEMS04 | HTVGNN-wc | 17.49 | 28.67 | 11.55 | 18.59 | 30.49 | 12.19 | 20.17 | 32.96 | 13.29 |
| HTVGNN | 17.38 | 28.53 | 11.47 | 18.18 | 29.86 | 11.93 | 19.41 | 31.68 | 12.74 | |
| PEMS08 | HTVGNN-wc | 12.90 | 21.34 | 8.45 | 13.96 | 23.35 | 9.13 | 15.53 | 25.83 | 10.06 |
| HTVGNN | 12.59 | 21.10 | 8.16 | 13.44 | 22.87 | 8.73 | 14.73 | 25.03 | 9.62 | |
Transformer-o: Transformer with temporal multi-head attention.
Transformer-tv: A Transformer that calculates attention scores with our proposed spatiotemporal multi-head attention.
The results of the two variants with different numbers of layers are presented in Table 5. It is evident that Transformer-tv outperforms Transformer-o significantly when the number of layers is 1 and 4. Our proposed time-varying mask matrix effectively mitigates the issue of inaccurate calculation in multi-head self-attention, while greatly enhancing temporal perception capability.
V-F2 Effect analysis of different types of graph structure in TVGCRM and HTVGNN
We have developed four variations of TVGCRM to underscore the significance of diverse manifestations of time-varying graphs, which are enumerated as follows:
TVGCRM: This is a coupled time-varying graph convolutional network.
TVGCRM-sl:Let different time slices be learned with different embeddings, but the spatial correlations learned in different time slices are not coupled.
TVGCRM-ag:: All time slices are learned using the same embedding, in which case the time-varying graph is the same as AGCRN.
TVGCRM-s:Replace coupled time-varying graphs with topological graphs.
The results presented in Table 6 unequivocally demonstrate the superiority of TVGCRM and its variants, namely TVGCRM-sl and TVGCRM-ag, over TVGCRM-s. This highlights the limitations of predefined graphs in accurately representing spatial correlations on traffic networks. Furthermore, the superior performance of TVGCRM-sl compared to TVGCRM-ag suggests that spatial correlations vary across different time steps, indicating that employing a single time-varying graph for learning is suboptimal. Moreover, the superiority of TVGCRM over TVGCRM-sl implies that considering the corresponding spatial correlations for different time steps is crucial as it allows capturing relevant information from other time steps to enhance modeling accuracy in current time steps.

### v3

@@ IV-C1 Graph learning
In order to capture the static spatial correlations between traffic nodes, we use two learnable embedding vectors to adaptively learn the static adjacency matrix at each time step. The static adjacency matrix of the time step is defined as Eq.(11):
where is a trainable node embedding, is a trainable bias of the node embedding . Herein, in order to ensure the diversity of at different time step, we set different bias for different time step during the training process, is used as the activation function.
In order to model the similarity of traffic flow between different traffic nodes, a traffic pattern graph is defined by calculating the correlations of historical traffic flows using the DTW algorithm [32]. In general, traffic pattern graph reflects a long-term traffic pattern. However, short-term traffic pattern may play an more important role in short-term traffic flow prediction. Therefore, in order to obtain more spatial topology-related information at different time steps, we have proposed a static graph learning module to improve the learning ability of the static adjacency matrix. At each time step, we can construct a static graph with the static adjacency matrix defined in Eq.(11). Then, in order to more accurately model static spatial correlations between different traffic nodes, we can further calculate the coupling time-varing adjacency matrix by the following Eq.(12):
where are learnable parameter matrices.after completing the graph coupling process,The coupling graph is utilized as the ultimate static adjacency matrix for spatial correlation modeling at each time step.
@@
where is the value of the hidden embedding at time step . is the mapped traffic features including short-term dynamic spatial traffic patterns implicit in current traffic features . is the node in . represents a splicing operation. indicates the dynamic traffic similarity between node and node in the road network.Then, using for normalization, the generation process of is as Eq.(15):
where is the value of at node and node , is the value of at time step .
Then,we think dynamic changes of spatial correlations occuring in the current time should belong to the actual existence and most active nodes. such as nodes which located in adjency matrix and DTW matrix.thus,we combine the adjency matrix and DTW matrix as the mask matrix of dynamic spatial correlations.Finally the dynamic adjacency matrix can be define as Eq.(16):
where is the value of at time step , is the value of at time step .
IV-C2 Coupling time-varying graph convolution gated recurrent unit
GRU is a variant of RNN which maintains the effectiveness of RNN in capturing temporal features while simplifying the parameters of the model, making it widely used in capturing temporal correlation. Based on the previous studies [33] [34], a coupling time-varying graph convolution gated recurrent unit (CTVGCRU) is constructed by replacing the matrix multiplication in GRU with a combination of graph convolutional network and an node adaptive parameter learning module. The specific form can be defined as Eq.(17):
where and are the input hidden embedding and output hidden embedding at the time step , respectively. represents the concatenation operation, and represent reset gate and update gate the time step , respectively. and represent the static graph operation and dynamic graph operation respectively., and , and , are learnable parameters in CTVGCRM.
| Dataset | PEMSD3 | PEMSD4 | PEMSD7 | PEMSD8 |
| # Nodes | 358 | 307 | 883 | 170 |
| # Edges | 866 | 340 | 340 | 277 |
| Time steps | 26208 | 16992 | 28224 | 17856 |
| Time span | 2018/9/1–2018/11/30 | 2018/1/1–2018/2/28 | 2017/5/1–2017/8/31 | 2016/7/1–2016/8/31 |
| Missing ratio | 0.672% | 3.182% | 0.452% | 0.696% |
| Time interval | 5 min | |||
| daily range | 0:00-24:00 | |||
@@ V-A Datasets
We validate the predictive performance of HTVGNN by conducting experiments on four real-world traffic datasets, namely PeMSD3, PeMSD4, PeMSD7, and PeMSD8. These datasets were collected by the California Department of Motor Transportation Performance Measurement System (PeMS). Please refer to Table I for detailed information.
@@ V-B Evaluation Metrics
In this study, we employed three effectiveness measures, namely Mean Absolute Error (MAE), Mean Absolute Percentage Error (MAPE), and Root Mean Square Error (RMSE), to assess the accuracy of the HTVGNN model as described in Eq.(18) to (20).
where and represent the predicted value and real value of traffic flow at node respectively.
@@ V-C Baselines
We compare the HTVGNN model against thirteen state-of-the-art baselines as outlined below:
@@
[37] ASTGCN introduces attention mechanisms into CNN and GCN to respectively model spatial and temporal dynamics of traffic data.
[33] AGCRN utilizes a node adaptive parameter learning to enhance GCN and combines it with GRU to capture spatial and temporal features of traffic data.
[38] GMAN uses spatial and temporal multi-attention mechanisms to model dynamic spatio-temporal correlations, and a gating mechanism is designed to better fuse spatio-temporal features.
@@
[40] a novel dual dynamic GCN traffic prediction method. DDSTGCN transforms the edges of traffic flow graph to its dual hypergraph and exploits their spatial and temporal property by GCN on hypergraph. [41] The encoder-decoder structure consists of a temporal Transformer and a spatial Transformer. The difference is that a recall module is added to the decoder to provide supplementary information for the prediction task. The designed DHM module is used to dynamically adjust the complexity of the model according to the complexity of the prediction task. [42] STWave decomposes complex traffic data into stable trends and fluctuating events, then models trends and events separately using dual-channel spatio-temporal networks. Finally, by fusing trends and events, we can predict reasonable future traffic flow. Additionally, we incorporate a novel query sampling strategy and graph wavelet-based graph position encoding into graph attention networks to efficiently and effectively model dynamic spatial correlations.
| Data | Encoder layers | Decoder layers | Batchsize | Heads | D | E | CTVGCRM layers | |
| PEMS03 | 1 | 1 | 16 | 8 | 64 | 8 | 8 | 2 |
| PEMS04 | 1 | 1 | 4 | 8 | 64 | 6 | 18 | 2 |
| PEMS07 | 1 | 1 | 8 | 8 | 64 | 10 | 24 | 2 |
| PEMS08 | 1 | 1 | 16 | 8 | 64 | 5 | 15 | 2 |
@@ V-D1 Dataset Preprocessing
Following recent methodologies [43], [44] and [45], we partition the training, validation, and test sets in a 6:2:2 ratio. To ensure more stable training, we employ z-score standardization. Additionally, we utilize historical traffic flow speed data from the past time slices (60 min) to predict future traffic flow speed in time slices (60 min). The input data is normalized using Z-Score as shown in equation (21).
@@ V-E Performance Comparison
Tables II and III present the comparison results of 12 prediction models on four datasets. Overall, our HTVGNN achieved exceptional prediction accuracy across all 12 prediction strides. The deep learning-based approaches outperformed traditional statistical-based approaches (VAR), indicating their efficacy in modeling highly nonlinear traffic flows. As depicted in Tables II and III, these approaches exhibited suboptimal prediction performance. Moreover, methods that capture spatiotemporal dependencies generally surpassed those solely modeling temporal dependencies (LSTM), underscoring the significance of spatial dependencies. DCRNN and STGCN employ predefined graphs to model spatial dependencies in road networks; however, their performance heavily relies on the quality of these predefined graphs. AGCRN utilizes adaptive graphs to capture spatial dependencies, which typically enhances predictions compared to predefined graphs. Nevertheless, both predefined and adaptive graphs remain static during the prediction stage, disregarding dynamic changes in the road network. In contrast, models that dynamically model spatial dependencies (GMAN, PDFormer) significantly outperform other approaches based on static graph structures (DCRNN, STGCN, GWN). This can be attributed to their ability to learn dynamically changing graph structures and provide a broader representation space for modeling more complex spatiotemporal dependencies.Compared to other models, HTVGNN incorporates a time-varying mask matrix to enhance the multi-head attention mechanism for accurately capturing dynamic temporal correlations. Moreover, HTVGNN employs a dedicated coupled graph learning method tailored for multi-head attention, effectively improving its performance in long-term prediction tasks. Additionally, to better model dynamic spatial correlations in road networks, we define a mask based on topological and semantic matrices to optimize the dynamic time-varying graphs. Overall, our model outperforms existing state-of-the-art approaches on four datasets, achieving superior prediction accuracy.
Additionally, in order to demonstrate the overall superiority of our model, we provide visual representations for each prediction point of HTVGNN and comparative models as depicted in Figure 4. In terms of the PEMS03 and PEMS04 datasets, our model exhibits the smallest cumulative error as the increase of the length of time steps, highlighting a distinct advantage of HTVGNN during long-term prediction stages. Furthermore, for PEMS07 and PEMS08 datasets, our model consistently maintains a significant advantage across all time steps.
V-F Abliation study
@@
: CTVGCRM should be utilized exclusively, with the exclusion of the enhanced temporal perception multi-head self-attention mechanism. This variant primarily demonstrates the function of the enhanced temporal perception multi-head self-attention.
@@
: Replace the time-varying graph in CTVGCRM with a topological graph, emphasizing its pivotal role in capturing temporal dynamic correlations.
@@
: Replace CTVGCRM with a topological graph-based graph convolution, which primarily emphasizes the role of the RNN module in HTVGNN. At this juncture, the number of layers remains unchanged.
@@
The experiments were conducted multiple times, and the optimal MAE, RMSE, and MAPE values for each variant on PEMSD4 and PEMSD8 are presented in Table V.
| PEMS04 | HTVGNN | w/o ETPMSA | w/o TV | w/o TR |
| MAE | 17.99 | 18.75 | 19.35 | 18.48 |
| RMSE | 29.74 | 30.55 | 31.29 | 30.08 |
| MAPE | 11.90 | 12.69 | 12.88 | 12.39 |
| PEMS08 | HTVGCN | w/o ETPMSA | w/o TV | w/o TR |
| MAE | 13.24 | 14.58 | 15.74 | 13.99 |
| RMSE | 22.67 | 23.66 | 25.06 | 23.23 |
| MAPE | 8.63 | 9.40 | 12.88 | 9.07 |
| # Layer | Variants | PEMS04 | PEMS08 | ||||
| MAE | RMSE | MAPE(%) | MAE | RMSE | MAPE(%) | ||
| L=1 | Transformer-o | 19.52 | 31.58 | 13.38 | 15.5 | 25.16 | 12.05 |
| Transformer-tv | 18.89 | 30.52 | 12.76 | 14.80 | 23.98 | 9.54 | |
| L=4 | Transformer-o | 18.56 | 30.90 | 12.35 | 14.83 | 24.43 | 9.45 |
| Transformer-tv | 18.45 | 30.16 | 12.13 | 13.75 | 23.28 | 8.98 | |
@@
The results presented in Table V demonstrate the superior performance of HTVGNN with ETPMSA, highlighting the crucial role played by ETPMSA in HTVGNN. Furthermore, the inferior performance of w/o TV compared to HTVGNN indicates that capturing spatial correlations through time-varying graphs is essential. Additionally, w/o TR being inferior to HTVGNN reveals the positive contribution of RNN in capturing temporal correlations.
@@
Given that our model is a hybrid time-varying graph convolutional neural network, this tutorial aims to dissect the impact of time-varying mask embedding and time-varying embedding on its performance by dividing the model into modules, ultimately analyzing their influence on the model’s performance.
@@
V-F1 Effect analysis of time-varying mask embedding in enhanced temporal perception multi-head attention
In order to demonstrate the effectiveness of our proposed time-varying mask embedding in enhanced temporal perception multi-head attention, the Transformer model with temporal trend-aware multi-head self-attention reported in [47] is selected as the comparison model and denoted as . For ease of description, the enhanced temporal perception multi-head attention with time-varying mask embedding is denoted as . Firstly, we investigate the impact of dimension of the mask embedding , and in the enhanced temporal perception multi-head self-attention. For the ease of analysis, we set the number of layers of and that of equal to . Then, we can obtain the optimal mask embedding. Figure 5 illustrates how different time-varing mask embeddings affect PEMS04 and PEMS08 datasets . As depicted in Figure 5, the optimal mask embeddings for PEMS04 and PEMS08 are respectively and .
@@ V-F1 Effect analysis of time-varying mask embedding in enhanced temporal perception multi-head attention
To highlight the advantages of enhanced temporal perception in multi-head self-attention, we set the number of layers to based on optimal mask embedding.
| Dataset | Graphs structures | MAE | RMSE | MAPE(%) |
| PEMS04 | CTVGCRM | 18.76 | 30.55 | 12.69 |
| CTVGCRM-sl | 18.93 | 31.05 | 12.73 | |
| CTVGCRM-ag | 19.35 | 31.29 | 12.88 | |
| CTVGCRM-s | 19.56 | 31.53 | 13.15 | |
| PEMS08 | CTVGCRM | 14.58 | 23.66 | 9.40 |
| CTVGCRM-sl | 15.08 | 24.25 | 9.77 | |
| CTVGCRM-ag | 15.24 | 24.38 | 10.66 | |
| CTVGCRM-s | 15.74 | 25.06 | 9.87 |
| PEMS04 | HTVGNN | w/o-tm | w/o-cg | w/o-bc |
| MAE | 17.99 | 18.53 | 18.35 | 18.53 |
| RMSE | 29.74 | 30.33 | 30.30 | 30.52 |
| MAPE(%) | 11.90 | 12.30 | 12.15 | 12.45 |
| PEMS08 | HTVGNN | w/o-tm | w/o-cg | w/o-bc |
| MAE | 13.24 | 14.10 | 13.79 | 14.55 |
| RMSE | 22.67 | 23.42 | 23.11 | 23.75 |
| MAPE(%) | 8.74 | 9.10 | 9.02 | 9.30 |
@@
| PEMS04 | w/o-cg | 17.49 | 28.67 | 11.55 | 18.59 | 30.49 | 12.19 | 20.17 | 32.96 | 13.29 |
| HTVGNN | 17.25 | 28.48 | 11.44 | 18.01 | 29.81 | 11.89 | 19.20 | 31.62 | 12.72 | |
| PEMS08 | w/o-cg | 12.90 | 21.34 | 8.45 | 13.96 | 23.35 | 9.13 | 15.53 | 25.83 | 10.06 |
| HTVGNN | 12.41 | 21.06 | 8.09 | 13.28 | 22.83 | 8.65 | 14.48 | 24.96 | 9.44 | |
The results of the experimental with different numbers of layers are presented in Table VI. It is evident that outperforms significantly when the number of layers is equal to and . Our proposed time-varying mask matrix effectively mitigates the issue of inaccurate calculation in multi-head self-attention, while greatly enhancing temporal perception capability.
V-F2 Effect analysis of different types of graph structure in CTVGCRM and HTVGNN
Similar to the preceding section, in order to evaluate the impact of diverse graph structures on CTVGCRM, we formulated four iterations of CTVGCRM as delineated follows:
: This is a coupling time-varying graph convolution gated recurrent network.
:Let different time slices be learned with different embeddings, but the spatial correlations learned in different time slices are not coupled.
: All time slices are learned using the same embedding, in which case the time-varying graph is the same as AGCRN.
:Replace coupled time-varying graphs with topological graphs.
The results presented in Table VII unequivocally demonstrate the superiority of CTVGCRM over its variants, namely CTVGCRM-sl ,CTVGCRM-ag and CTVGCRM-s.This highlights the limitations of predefined graphs in accurately representing spatial correlations on traffic networks. Furthermore, the superior performance of CTVGCRM-sl compared to CTVGCRM-ag suggests that spatial correlations vary across different time steps, indicating that employing a single time-varying graph for learning is suboptimal. Moreover, the superiority of CTVGCRM over CTVGCRM-sl implies that considering the corresponding spatial correlations for different time steps is crucial as it allows capturing relevant information from other time steps to enhance the modeling accuracy of the current time step.